# Bidirectional Decoding: Improving Action Chunking via Closed-Loop Resampling

### v1

@@ 5.5.1 Dynamic Placing
Following previous works [4, 13], we use a Franka Panda as the robot hardware and the vision-based diffusion policy for its operation. The robot is equipped with two cameras: one ego-centric camera mounted at the wrist of the robot, one third-person camera mounted at a static bracket. Both cameras provide visual observations at a resolution of $256\times 256$ pixels. The robot operates at a frequency of 10 Hz, with a prediction horizon of 16 time steps.
@@
5.5.2 Dynamic Picking
Task.
Policy and Robot.
We utilized the publicly available diffusion policy checkpoint from UMI [22] without any additional fine-tuning. Notably, the policy was originally trained using demonstrations in a static setting, where the cup’s position remained constant throughout the task. Our experimental setup mirrored the one described by UMI, using the same UR5 robot hardware. This allowed us to directly evaluate the policy’s transferability to a dynamic environment, where the cup’s position changes during the task.
Evaluation.
Results.
@@
Other experiments.
Please refer to Appendix A for additional analyses and ablations.
6 Conclusion
Acknowledgments
Appendix A Additional Experiments
A.1 One-dimensional Simulations
Action Horizon
Noise Level 1 3 5 7 10
0.0 4.21 1.75 1.55 1.28 0.00
0.4 0.55 0.30 0.95 0.53 0.93
0.8 0.04 0.98 1.23 1.26 1.44
A.2 Other Horizon Choices
Setup. Our work builds on the premise that the prediction horizon is longer than the context length, as commonly designed for recent policies. While BID mitigates the inherent limitations of this design choice through test-time decoding, an important question remains: could extending the history context itself yield stronger policies? To understand this, we trained diffusion policies with varying combinations of prediction horizons and context lengths on the Push-T task. Specifically, we use a short context length ($c=2$) and a short prediction horizon ($h=2$) as our baseline, and incrementally increase these parameters to larger values ${6,10,14}$ to assess their impact.
Result. Fig. 12 compares the performance of the policy learned with different $\Delta h=h-c$. As expected, the policy with both a short prediction horizon and a short context length struggles to capture long-range temporal dependencies, leading to suboptimal performance. Interestingly, extending the context length initially boosts performance ($\Delta h=-4$), but this trend reverses as the context length becomes too long ($\Delta h\leq-8$), likely due to overfitting to an increased number of spurious features. In contrast, expanding the prediction horizon results in more robust performance improvements, validating its pivotal role in policy design given limited demonstrations.
A.3 Ablation Study of Forward Contrast
Setup.
To understand the effect of forward contrast (Equation 8), we evaluate the full version of our method against three reduced variants: without forward contrast, without positive samples (negative samples only), and without negative samples (positive samples only). Similar to Sec. 5.3, our ablation study is conducted in the representative Push-T task.
Result.
Fig. 14 summarizes the result of this ablation study. Notably, both positive and negative samples are essential for effective sample selection, and omitting either leads to significant performance declines. We conjecture that, without negative samples, our decoding method reduces to an approximate maximum a posteriori estimation, which can result in suboptimal decisions due to modeling errors. Conversely, without positive samples, the selected samples may be biased towards rare instances. This result highlights the importance of both components and suggests the potential for extending this paradigm in future work.
Figure 13: Effect of positive and negative samples on forward contrast. We measure the improvements over the vanilla baseline in the Push-T task, relative to the full version of BID. Using only positive or only negative samples does not achieve the full performance gains seen with the full contrastive.
Figure 14: Effect of the decay rate for the exponential moving average. In each task, we measure the relative performance among different decay rates. The optimal decay rate varies by task, leading to a practical challenge of identifying a universal temporal ensembling strategy [3].
A.4 Challenges for Temporal Ensebmling
EMA exhibits competitive performance in Table 1. However, tuning its decay rate can be difficult in practice. Fig. 14 shows the sensitivity of EMA to the decay rate across three different tasks, where the optimal choices differ significantly. We conjecture that this high sensitivity stems from the variability in the latent strategies between consecutive predictions. When consecutive predictions follow similar strategies, a lower decay rate (i.e., stronger moving average) can enhance smoothness and improve performance. Conversely, when consecutive predictions diverge in their underlying strategies, averaging them can introduce adverse effects. Our method promotes coherence in latent strategies and thus effectively complements temporal ensembling, as evidenced in Fig. 6.
Appendix B Additional Details
Simulation Details.
Push-T: We adopt the Push-T environment introduced in [4], where the goal is to push a T-shaped block on a table to a target position. The action space is two-dimensional end-effector velocity control. The training dataset contains 206 demonstrations collected by humans.
Robomimic: We use five tasks in the Robomimic suite [56], namely Lift, Can, Square, Transport, and Tool Hang. The training dataset for each task contains 300 episodes collected from multi-human (MH) demonstrations.
Franka Kitchen: We use the Franka Kitchen environment from [57], featuring a Franka Panda arm with a seven-dimensional action space and 566 human-collected demonstrations. The learned policy is evaluated on test cases involving four or more objects (p4), a challenging yet practical task for robotic manipulation in household contexts.
Implementation Details.
Our implementations are built upon the official code of Diffusion Policy [4], with modifications made solely to the inference process. Our simulation experiments use state inputs, while our real-world experiments utilize visual inputs. For each simulation task, we use the best checkpoint available from https://diffusion-policy.cs.columbia.edu/data/experiments and evaluate it in the closed-loop operation, i.e., action horizon is set to 1. For forward contrast, we train the weak policy for 50-100 epochs, resulting in a diverse set of suboptimal plans. The core hyperparameters are summarized in Table 5.
For VQ-BeT [6], we use the best public checkpoint from the LeRobot Repository [58] as the strong policy whereas we used a checkpoint trained for 100000 iterations as the weak policy. Backward coherence, forward contrast and BID require sample diversity, so we select temperature=0.5 for these methods whereas for the vanilla sampling, we used the default value temperature=0.1. Our code will be released publicly.
name value
batch size $N$ 30
mode size $K$ 10
prediction horizon $l$ 16
temporal coherence decay $\rho$ 0.9
moving average decay $\lambda$ 0.75
Table 5: Default hyper-parameters in our experiments.

### v2

@@ B.2.1 Dynamic Placing
Following previous works (Chi et al., 2023; Prasad et al., 2024), we use a Franka Panda as the robot hardware and the vision-based diffusion policy for its operation. The robot is equipped with two cameras: one ego-centric camera mounted at the wrist of the robot, one third-person camera mounted at a static bracket. Both cameras provide visual observations at a resolution of $256\times 256$ pixels. The robot operates at a frequency of 10 Hz, with a prediction horizon of 16 time steps.
@@
B.2.2 Dynamic Picking
Task.
Policy.
We utilized the publicly available diffusion policy checkpoint from UMI (Chi et al., 2024) without any additional fine-tuning. Notably, the policy was originally trained using demonstrations in a static setting, where the cup’s position remained constant throughout the task. Our experimental setup mirrored the one described by UMI, using the same UR5 robot hardware. This allowed us to directly evaluate the policy’s transferability to a dynamic environment, where the cup’s position changes during the task. Due to the absence of an early checkpoint, we omitted negative samples in forward contrast, focusing solely on positive consistency discussed in Sec. A.3.
Evaluation.
Results.